OpenStudy (anonymous):

Find approximate root x^2-8x-12 between x=9 and x=10

OpenStudy (amistre64):

simple enough with the quadratic formula

OpenStudy (anonymous):

Is it alright if you show me step by step on how to do it and what numbers go where? please.

OpenStudy (amistre64):

when given the structure:\[ax^2+bx+c\]we can find the roots simple enough by using the formula:\[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]or by completing the square

OpenStudy (anonymous):

So would my answer be -2,6. I did completing the square, but I dont understand when they mean between 9 and 10

OpenStudy (amistre64):

most likely there are 2 roots and they want the one that is between 9 and 10

OpenStudy (amistre64):

x^2-8x-12 ; 8/2 = 4; 4^2 = 16 x^2-8x+16 -16 -12 = 0 (x-4)^2 -28 = 0 x-4 = sqrt(28) x = 4 + sqrt(28)

OpenStudy (amistre64):

28 is close to 25 sqrt(25) = 5 so this is a bit bigger than 9 overall

OpenStudy (anonymous):

I got 9.29. Is that correct

OpenStudy (amistre64):

my calculator says the same thing

OpenStudy (anonymous):

okay thank you so much.

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

you helped me a lot and I really appreciate it.